An electrical firm manufactures light bulbs that have a length of life that is approximately normally distributed, with mean equal to 801 hours and a standard deviation of 98 hours. Find the probability that a random samples of 19 bulbs will have an average life between 782 and 818 hours.


1
Expert's answer
2022-02-15T13:51:36-0500

"\mu=801\\\sigma=98\\n=19"

Let the random variable "X" represent the length of life of light bulbs. We are required to find the probability, "p(782\lt \bar x\lt 818)=p({782-\mu\over{\sigma\over\sqrt{n}}}\lt {\bar x-\mu\over{\sigma\over\sqrt{n}}}\lt{818-\mu\over{\sigma\over\sqrt{n}}})=p({782-801\over{98\over\sqrt{19}}}\lt Z\lt {{818-801\over{98\over\sqrt{19}}}})=p(-0.85\lt Z\lt0.76)=\phi(0.76)-\phi(-0.85)=0.7764-0.1977=0.5787"

Therefore, the probability that a random sample of 19 bulbs will have an average life between 782 and 818 hours is 0.5787